OpenStudy (anonymous):

Solve for x. Identify if extraneous or not. Show work. -4(sqrtx+9)=20 Please help!! I am so lost :(

OpenStudy (anonymous):

is it \[-4\sqrt{x+9}=20\]?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

I don't understand... How can you go from that whacky 5 letter variable to just x?

OpenStudy (anonymous):

sqrt is short for square root

OpenStudy (anonymous):

oh lol, fair enough

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

Help me please I really don't get imaginary numbers

OpenStudy (anonymous):

sec

OpenStudy (anonymous):

hmmmm, well to be totally honest, I haven't exactly learnt imaginary numbers yet lol. I am getting x=sqrt -8

OpenStudy (anonymous):

is that similar to your answer so far?

OpenStudy (anonymous):

I was getting like x=-5i but that 9 is messing me up

OpenStudy (anonymous):

mm another way that I can work it out results in x=sqrt-5 -9

OpenStudy (anonymous):

Yeah but I need an imaginary number (i) to get the negative four to turn into a positive number when multiplied by the radical

OpenStudy (tkhunny):

What are you doing? \(-4\sqrt{x+9} = 20\) has no solution at all, Real, Complex, or otherwise. This is obtained by inspection. By definition, \(\sqrt{x+9} \ge 0\) -- Done!